Yesterday, Mar 12 2019 20:48, took 2 minutes of data with arm and MICH flashes and OMCs close to resonance to have light arriving on the new B1 PDs. This allowed to measure the transfer function between the Audio and DC channel and design blending filters. The Audio channel low-frequency cut-off is measured to be a double pole at 1.8Hz for PD1 and at 1.75Hz for PD2.
In addition, in LAPP we have used the spare PCAL laser to measure the absolute response of the DC channel and the Audio channel. As expected, with the new preamplifier the responses are no longer flat due to the poorer DC/Audio separation at low frequency and Audio/RF separation at high frequency. We have added corrective shaping in the step where the DC/Audio data is downsampled to 20kHz (before blending). For the low frequency, a zero at 1.4Hz and pole at 1.6Hz has been added to both DC and Audio channels, as the blending filter was measured without this correction to either channel. At high frequency we have added a zero 16kHz to the Audio channel.
Figure 1 shows the measured electronic noise of the blended signal calibrated into mW using data 3min of data from 16:09 UTC, and a by hand fit that will be used in the noise budget.